# Flight Demonstration and Model Validation of a Prototype Variable-Altitude Venus Aerobot

### v1

@@ 3.2.1 Dynamics Model
The horizontal motion is similarly governed (with buoyancy and gravity neglected) and all masses lumped as points as the multibody swinging dynamics were not the focus of this experiment. Note that for this analysis, we itemize the total mass including helium gas as:
@@ 3.2.2 Shape Model
The balloon shape varies with time, and accordingly must be computed each timestep to determine the vertical drag area. The FLOATS shape computation assumes that the balloon shape is axisymmetric, is in a shape equilibrium, and is parameterized in terms of arc length. The shape takes a two-lobed path from the bottom apex to top apex of the ZP balloon, pushed towards and away from the central balloon axis as dictated by the circumferential stress on the ZP balloon material from the internal helium gas and the SP balloon inside. The arc length has three partitions:
@@
    The base section, where the ZP balloon shape lies against the SP balloon envelope. Here the ZP circumferential stress is negative, compressing the two envelopes together. The initial meridional stress from the gondola weight and net buoyancy from the internal SP balloon is applied to this section (see Appendix 6.3).
@@
    The mid section, where the ZP balloon shape is loose and the circumferential stress in the ZP envelope is exactly or very nearly zero, as the envelope is in equilibrium.
@@
    The apex section, where the ZP envelope is taut. Here the ZP circumferential stress is positive, pushing the ZP balloon shape to its maximum extent. Physically, this happens because the helium gas inside is buoyant, and therefore collects at the top. As the ZP balloon envelope has high elastic modulus it does not stretch enough to meaningfully affect the shape, so when the balloon material is pulled taught it will very nearly follow the as-fabricated fully inflated geometry.
@@
This mid section of the balloon shape (Fig. 8b) is governed by the following set of ordinary differential equations in terms of the arc length $s$ for four variables - the shape angle, normalized meridional stress, radius from central axis, and height:
@@
$\theta(s_{0})$ $=\frac{\pi}{2}-\beta$
@@
$\theta(s_{\ell})$ $=\theta_{\text{infl}}(s_{\ell})$
@@
A fundamental limitation of this model is that it ignores the excess slack (i.e. ruffles of material) around the ZP balloon circumference, instead assuming the shape is perfectly axisymmetric with a gas bubble diameter that varies with height. While our model is therefore sufficient for the gross estimation of projected areas, these ruffles can add drag and asymmetries in film loading. We accordingly design the aerobot to higher wind and tensile load than expected in flight, and use field data to tune and validate the performance.
@@ 3.2.3 Aerodynamic Model
From the governing equation Eq. 3.2.1, the primary aerodynamic forces exerted on the system are the drag and virtual mass, expressed as detailed by Carlson and Horn [25]. Notably, while these forces are significant for the balloon-in-a-balloon configuration, they are negligible for the payload given the vast difference in volume and drag area.
@@
Choosing a $C_{D_{\text{top}}}$ and $C_{D_{\text{side}}}$ is difficult for balloons, as the flexible outer envelope generally does not support re-attachment of a turbulent boundary layer (unlike a sphere) [25]. Furthermore, the "Morison Equation" employed in Eq. 3.2.1 for the drag and virtual mass is effective for steady-state forces and initial transient forces, but less so for the intermediate-time forces [26]. As such, the virtual mass and drag coefficients need to be tuned to experimental data at the appropriate movement timescale [27]. We choose $C_{D_{\text{top}}}=0.8$ and $C_{D_{\text{side}}}=1$ given good agreement with the indoor calibration flight data in Izraelevitz et al. [13], a $C_{m}=0.2$ from ellipsoidal data in Hoerner [26], and leave higher fidelity fluid-structure interaction simulations to future work.
@@ 3.2.4 Thermodynamics Model
[…]
@@
Some simplification of the gas temperature in Eqs. 17 and 15 is possible using $c_{p}=c_{v}+[\text{gas constant}]$ as in other balloon thermodynamics derivations [25, 28], but keeping everything explicit allows for full enumeration of the enthalpy paths.
@@ 3.2.5 Heat Transfer Model
The heat transfer model considers convection and radiation as the primary forms of heat transfer for the gases and envelopes. A full description of heat flows $q_{i}$, $q_{SP}$, $q_{ZP}$ into nodes is described in Patel and Cameron [29], but we include an abbreviated version here for context.
The SP and ZP heat flows from their respective envelope boundaries $i=1,2,3,4$ in Fig. 9b to their internal helium gas are given as:
@@
The heat flows to each of the four envelope nodes $i=1,2,3,4$ are then given by convective and radiative (both solar and infrared) terms:
@@
$\dot{q}_{\text{3}}$ $=\dot{q}_{\text{solar, 3, sidewelling}}+\left(\dot{q}_{\text{infrared, 3, sidewelling}}-\dot{q}_{\text{infrared, }3,\text{out}}\right)+\sum_{i=\text{atm, ZP}}\dot{q}_{\text{convection, }i\rightarrow 3}+\sum_{i=2,4,\text{ZP}}\dot{q}_{\text{infrared, }i\rightarrow 3}$
$\dot{q}_{\text{4}}$ $=\dot{q}_{\text{solar, 4, downwelling}}+\left(\dot{q}_{\text{infrared, 4, downwelling}}-\dot{q}_{\text{infrared, }4,\text{out}}\right)+\sum_{i=\text{atm, ZP}}\dot{q}_{\text{convection, }i\rightarrow 4}+\sum_{i=2,3,\text{ZP}}\dot{q}_{\text{infrared, }i\rightarrow 4}$ (23)
The fluxes impinging on the external envelope of the aerobot are broken into upwelling, sidewelling, and downwelling irradiances $E$ of the environment:
$\dot{q}_{\text{solar, i, up/down/sidewelling}}$ $=\alpha_{i}A_{i}E_{\text{up/down/sidewelling}}$
$\dot{q}_{\text{infrared, i, up/down/sidewelling}}$ $=\epsilon_{i}A_{i}E_{\text{up/down/sidewelling}}$
@@
$\dot{q}_{\text{infrared, }i\rightarrow j}=\sigma\left(T^{4}_{i}-T^{4}_{j}\right)\left/\left(\frac{1-\epsilon_{i}}{\epsilon_{i}A_{i}}+\frac{1}{f_{i\rightarrow j}A_{i}}+\frac{1-\epsilon_{j}}{\epsilon_{j}A_{j}}\right)\right/$ (25)
@@
$\dot{q}_{i=\text{convection, atm }\rightarrow j}$ $=\frac{\text{Nu}_{\text{forced}}\cdot k_{\text{atm}}}{l_{\text{atm}\rightarrow j}}A_{\text{atm}\rightarrow j}(T_{i}-T_{j})$
@@
$\text{Nu}_{\text{forced}}$ =$=\left\{\begin{matrix}0.37\text{Re}^{0.6}&\text{Re}\leq 5.38\cdot 10^{5}\\ 0.74\text{Re}^{0.6}&\text{otherwise}\end{matrix}\right\}$ (26)
@@ 3.2.6 Atmospheric model
Figure 10: Atmospheric pressure, temperature, and winds recorded on (a, c) Flight 1 and (b, d) Flight 2. Wind velocity was computed in the inertial reference frame in the “up" direction. These values were computed by vectorially adding the vertical velocity of the balloon with the wind velocity recorded on the anemometer on the gondola.
@@
In post-flight analysis, the multi-sensor dataset collected in the outdoor free flights was used to overwrite EarthGram as new inputs to the FLOATS model to account for the exact weather of the day. Figures 10 and 11 show as-measured atmospheric property inputs such as temperature, pressure, winds, and solar radiation supplied to the FLOATS model. Data used within the FLOATS model were corrected based on post-flight calibration as discussed in Appendix Appendix. Venting and pumping times were supplied to the FLOATS model based on data collected on board as well.
Atmospheric temperature was measured via two methods: (1) an ultrasonic anemometer sonic temperature and (2) a thermistor exposed to the surrounding air. The thermistor was observed to have a large response time lag and, in comparison, the sonic temperature showed more consistent values between ascent and descent albeit some noise. Accordingly, a filtered spline fit of the sonic temperature (Figure 10a-b, red) was used as temperature input to FLOATS.
The anemometer also provided 3-D winds measured on the aerobot gondola (Figure 10c-d, black). These were rotated into the inertial frame and vectorially added to the aerobot inertial velocity (obtained from the INS on the gondola) to generate inertial wind velocities, which were then used by FLOATS.
Lastly, up- and down-welling shortwave (visible solar) radiation and longwave (infrared) radiation were supplied to FLOATS as inputs (Figure 11). The radiation was assigned to one of three sources:
  1. 1.
    A planar ground source in both upwelling longwave (ground temperature) and upwelling shortwave (ground albedo). This source strength was determined from the pyranometers and pyrgeometers on the gondola.
  2. 2.
    A planar sky source in downwelling longwave (sky temperature). This source strength was determined from the ground station pyrgeometer. Accidental shadows cast on the groundstation by the field team were filtered out in post-processing. For both our flights, this sky temperature was cooler than the ground temperature.
  3. 3.
    A solar point source in direct shortwave (sunlight). This source strength was computed from all four pyranometers as none permanently face the morning sun. A sidewelling solar flux was first computed as an envelope function from the side-mounted pyranometers, accounting for the motion in and out of the sunlight as the aerobot rotates. Subsequently, the upwelling flux from the bright ground was subtracted from this sideways signal with a viewfactor of 1/2, and then direct solar magnitude was determined from the hypotenuse of horizontal and vertical fluxes as:
    $E_{\text{direct, solar}}=\sqrt{\left(E_{\text{sidewelling, solar}}-\frac{1}{2}E_{\text{upwelling, solar}}\right)^{2}+E_{\text{downwelling, solar}}^{2}}$ (27)
    Given the early morning test flights with the sun close to the horizon, however, the hypotenuse effect had little influence. The resultant direct solar flux is nearly identical to the sidewelling signal, but the distinction would be important for future longer duration flights closer to local noon.
All inputs could be supplied to FLOATS as functions of either altitude or time, depending on the phenomenon. We chose to parameterize the measured shortwave radiation and winds by time, as solar incidence angle is primarily time-dependent at low altitude, and the vertical wind gusts proved to be non-recurring disturbances. We decided to parameterize the measured atmospheric pressure and temperature by altitude instead of time (with high confidence for pressure and medium confidence on temperature; see Results Section 4), as these properties showed repetition on the aerobot ascent and descent - presumably given the long timescale of the atmosphere change in the morning compared to the balloon motion. Longwave radiation was nearly constant, but showed minor recurrence between ascent and descent and therefore was also parameterized by altitude.
@@ 4 Results
The main output comparisons for performance evaluation of the FLOATS model (namely aerobot altitude, SP balloon pressure, and gas and skin temperature) are shown in Figures 13, 13, and 14.
@@
Good agreement is also seen between predicted and observed SP absolute pressure values in both flights (Fig. 13), with approximately 2 kPa accumulated error over time as the driving pressure for moving gas is dependent on the precise altitude. We also note good agreement (Fig. 14) between the simulated and observed gas and skin temperatures with the ZP gas temperature. The thermistor measurements at the top of the aerobot are the most consistent with FLOATS predictions. Typical thermal error between FLOATS predictions and temperature measurements is within ±$\pm 10^{\circ}$, with better agreement during Flight 1 compared to Flight 2. The sudden temperature rise in FLOATS after activating the termination poppet is likely due to the poppet enthalpy flow not being included in Eq. 18, as free-flight thermal behavior was the priority, so the ZP heat gets unrealistically distributed on the decreasing amount of gas and raises temperature.
Figure 12: Simulated and observed altitudes for both outdoor flights. (a) Flight 1 on July 14, 2022 and (b) Flight 2 on July 16, 2022. Green sections labeled “vent" in the flight data indicate times when vent was active, red sections labeled “pump" indicate times when pump was active. The label “termination" indicates time at which the poppet was activated to terminate the flight. We assumed 100 grams of free lift in the FLOATS simulation for Flight 1 to match the peak altitude measured given lack of reliable lift measurement for this first flight, but use the as-measured free lift of 320g for Flight 2.
Figure 13: Simulated and observed SP balloon pressure for both outdoor flights. (a) Flight 1 on July 14, 2022 and (b) Flight 2 on July 16, 2022. As before, green sections labeled “vent" in the flight data indicate times when vent was active, red sections labeled “pump" indicate times when pump was active. Pump noise is an artifact of compression waves near the pressure transducer from the oscillating pump pistons.
Figure 14: Simulated and observed temperature at various nodes during the outdoor flight. The dotted line shows FLOATS simulation predictions while the solid line shows observed data. Note that the T2 node and SP node were not measured, and T3 was measured in two locations on opposite sides of the ZP envelope (shown by black and gray solid lines). Corrections from the post-flight calibration of thermistors (see appendix) were applied.
In our analysis, we found that the heat transfer model within FLOATS is the largest source of sensitivity in the outputs. Notably, the simulated aerobot altitude showcases the strongest sensitivity to the atmospheric temperature near the surface and the solar radiation input. FLOATS cannot currently accommodate a non-steady atmosphere and only accepts the measured atmospheric variation with altitude – thus variations in temperature at the same altitude during the approximately hour-long flights of the aerobot were ignored. This is a key limitation of the model, especially in the morning as hot air near the ground begins to rise and mix, decreasing the air density. By contrast, FLOATS does simulate the rise in temperature of the ZP helium gas in sunlight. Accordingly, while FLOATS includes the increasing buoyancy effect from the warm helium it simultaneously ignores the decreasing buoyancy effect from the warm air.
[…]
Future FLOATS modeling work will focus on these thermal excursions and requires longer duration flights for validation under many hours of variable solar forcing and atmospheric temperature change.

### v2

@@ 3.2.1 Dynamics Model
Where the displaced volume V is computed from the sum of the fixed SP balloon volume and the variable ZP balloon volume given by the ideal gas law. The horizontal motion is similarly governed (with buoyancy and gravity neglected) and all masses lumped as points as the multibody swinging dynamics were not the focus of this experiment. Note that for this analysis, we itemize the total mass including helium gas as:
@@ 3.2.2 Shape Model
The balloon shape varies with time, and accordingly must be computed each timestep to determine the vertical drag area. The FLOATS shape computation assumes that the balloon shape is axisymmetric, is in a shape equilibrium, and is parameterized in terms of arc length. The shape takes a two-lobed path from the bottom apex to top apex of the ZP balloon (Fig. 8b), pushed towards and away from the central balloon axis as dictated by the circumferential stress on the ZP balloon material from the internal helium gas and the SP balloon inside. The arc length has three partitions:
@@
    The base section $(s<s_{0})$: Here the ZP balloon shape lies against the SP balloon envelope. Here the ZP circumferential stress is negative, compressing the two envelopes together. The initial meridional stress from the gondola weight and net buoyancy from the internal SP balloon is applied to this section (see Appendix 6.3).
@@
    The mid section $(s_{0}<s<s_{l})$: Here the ZP balloon shape is loose and the circumferential stress in the ZP envelope is exactly or very nearly zero, as the envelope is in equilibrium.
@@
    The apex section $(s>s_{l})$: Here the ZP envelope is taut. Here the ZP circumferential stress is positive, pushing the ZP balloon shape to its maximum extent. Physically, this happens because the helium gas inside is buoyant, and therefore collects at the top. As the ZP balloon envelope has high elastic modulus it does not stretch enough to meaningfully affect the shape, so when the balloon material is pulled taught it will very nearly follow the as-fabricated fully inflated geometry.
@@
This mid section of the balloon shape is governed by the following set of ordinary differential equations in terms of the arc length $s$ for four variables - the shape angle, normalized meridional stress, radius from central axis, and height:
@@
$\theta_{\text{ZP}}(s_{0})$ $=\frac{\pi}{2}-\beta$
@@
$\theta_{\text{ZP}}(s_{\ell})$ $=\theta_{\text{infl}}(s_{\ell})$
@@
A fundamental limitation of this model is that it ignores the excess slack (i.e. ruffles of material) around the ZP balloon circumference, instead assuming the shape is perfectly axisymmetric with a gas bubble diameter that varies with height. While our model is therefore sufficient for the gross estimation of projected areas, and matches well with photo outlines of the balloon shape, these ruffles can add drag and asymmetries in film loading. We accordingly design the aerobot to higher wind and tensile load than expected in flight with extra layers of Kapton near the endfittings to spread the point load, and use field data to tune and validate the performance.
@@ 3.2.3 Aerodynamic Model
From the governing equation Eq. 3.2.1, the primary aerodynamic forces exerted on the system are the drag and virtual mass, expressed in detail by Carlson and Horn [27]. Notably, while these forces are significant for the balloon-in-a-balloon configuration, they are negligible for the payload given the vast difference in volume and drag area.
@@
Choosing a $C_{D_{\text{top}}}$ and $C_{D_{\text{side}}}$ is difficult for balloons, as the flexible outer envelope generally does not support re-attachment of a turbulent boundary layer (unlike a sphere) [27]. Furthermore, the "Morison Equation" employed in Eq. 3.2.1 for the drag and virtual mass is effective for steady-state forces and initial transient forces, but less so for the intermediate-time forces [28]. As such, the virtual mass and drag coefficients need to be tuned to experimental data at the appropriate movement timescale [29]. We choose $C_{D_{\text{top}}}=0.8$ and $C_{D_{\text{side}}}=1$ given good agreement with the indoor calibration flight data in Izraelevitz et al. [13], a $C_{m}=0.2$ from ellipsoidal data in Hoerner [30], and leave higher fidelity fluid-structure interaction in computation fluid dynamics (CFD) simulations to future work on how these coefficients should change independent of reference areas.
@@ 3.2.4 Thermodynamics Model
Figure 9: Thermal nodes and control volumes used in thermodynamic modeling. (a) Work and gas enthalpies, along with the six thermal nodes ($i=1,2,3,4,$ ZP, SP). (b) Temperature sensor placement, with the exception of Node 2 and Node SP which were not measured due to access constraints. Node 3 was measured in two locations 3a and 3b on opposite sides. Photo courtesy NASA/JPL-Caltech, ©2022 California Institute of Technology.
@@
Some simplification of the gas temperature in Eqs. 17 and 15 is possible using $c_{p}=c_{v}+\overline{R}$ as in other balloon thermodynamics derivations [27, 31], but keeping everything explicit allows for full enumeration of the enthalpy paths.
@@ 3.2.5 Heat Transfer Model
The heat transfer model considers convection and radiation for all nodes (Fig. 9a). Accordingly, the lifting gas nodes (SP & ZP) see convective and radiative heating to the envelope, and the envelope nodes (i = $1,2,3,4$) see convective and radiative heating to both the lifting gas and the outside environment. Starting with the gas nodes, their total heat flux is given as:
@@
Note that these two gas nodes (SP & ZP) do not see the outside environmental fluxes directly. The heat flows to each of the four envelope nodes on their boundaries ($i=1,2,3,4$) are then given by convective and radiative (both solar and infrared) terms to the outside environment:
@@
[…]
The fluxes impinging on the external envelope of the aerobot are broken into upwelling, downwelling, and direct irradiances $E$ of the environment from three sources (illustrated in Fig. 9):
  1. 1.
    A planar ground source in both upwelling longwave $E_{\text{upwelling, infrared}}$ (ground temperature) and upwelling shortwave $E_{\text{upwelling, solar}}$ (ground albedo).
  2. 2.
    A planar sky source in downwelling longwave $E_{\text{downwelling, infrared}}$ (sky temperature).
  3. 3.
    A solar point source in direct shortwave $E_{\text{direct, solar}}$ (sunlight).
[…]
$\dot{q}_{\text{solar, 1, upwelling}}$ $=\alpha A_{1}E_{\text{upwelling, solar}}$
[…]
The infrared viewfactors assume diffuse upwards and downwards sky and ground sources as:
$\dot{q}_{\text{infrared, 1, upwelling}}$ $=\epsilon A_{1}E_{\text{upwelling, infrared}}$
$\dot{q}_{\text{infrared, 3, sidewelling}}$ $=\epsilon A_{3}\left(\frac{1}{2}E_{\text{upwelling, infrared}}+\frac{1}{2}E_{\text{downwelling, infrared}}\right)$
$\dot{q}_{\text{infrared, 4, downwelling}}$ $=\epsilon A_{4}E_{\text{downwelling, infrared}}$
@@
$\dot{q}_{\text{infrared, }i\rightarrow j}=\sigma\left(T^{4}_{i}-T^{4}_{j}\right)\left/\left(\frac{1-\epsilon_{i}}{\epsilon_{i}A_{i}}+\frac{1}{f_{i\rightarrow j}A_{i}}+\frac{1-\epsilon_{j}}{\epsilon_{j}A_{j}}\right)\right$. (26)
@@
$\dot{q}_{\text{convection, atm }\rightarrow j}$ $=\frac{\text{Nu}_{\text{forced}}\cdot k_{\text{atm}}}{l_{\text{atm}\rightarrow j}}A_{\text{atm}\rightarrow j}(T_{i}-T_{j})$
@@
$\text{Nu}_{\text{forced}}$ $=\left\{\begin{matrix}0.37\text{Re}^{0.6}&\text{Re}\leq 5.38\cdot 10^{5}\\ 0.74\text{Re}^{0.6}&\text{otherwise}\end{matrix}\right\}$ (27)
@@ 3.2.6 Atmospheric model
Figure 10: Atmospheric pressure, temperature, and inertial-frame winds recorded from (a, c) Flight 1 and (b, d) Flight 2. Wind velocity in the inertial frame were computed by vectorially adding the vertical velocity of the balloon with the wind velocity recorded on the anemometer on the gondola.
Figure 11: Solar radiation from pyranometers for (a) Flight 1 and (b) Flight 2. Longwave IR from pyrgeometers for (c) Flight 1 and (d) Flight 2. Upwelling and side radiation are measured on the gondola (150 and 180 FOV respectively), downwelling radiation from the ground station with 180 FOV.
In post-flight analysis, the multi-sensor dataset collected in the outdoor free flights was used to overwrite EarthGram as new inputs to the FLOATS model to account for the exact weather of the day. Figures 10 and 11 show as-measured atmospheric property inputs such as temperature, pressure, winds, and solar radiation supplied to the FLOATS model. Data used within the FLOATS model were corrected based on post-flight calibration as discussed in Appendix Appendix: Sensor Details and Shape Model Supplement. Venting and pumping times were supplied to the FLOATS model based on data collected on board as well.
Atmospheric temperature was measured via two methods: (1) an ultrasonic anemometer sonic temperature and (2) a temperature sensor exposed to the surrounding air. The temperature sensor was observed to have a large response time lag (rise time of approximately 4 minutes leading to errors of up to $5^{\circ}$C) and, in comparison, the sonic temperature showed more consistent values between ascent and descent albeit some noise. Accordingly, a filtered spline fit of the sonic temperature (Figure 10a-b, red) was used as temperature input to FLOATS.
The ultrasonic anemometer also provided 3-D winds measured on the aerobot gondola (Figure 10c-d, black). These were rotated into the inertial frame and vectorially added to the aerobot inertial velocity (obtained from the INS on the gondola) to generate inertial wind velocities, which were then used by FLOATS. Ground winds from the cup anemometer at the launch site were not logged, only used for launch go/no-go decisions.
Diffuse shortwave and longwave radiation were measured from a variety of locations (Figure 11) and then assigned to irradiance sources for FLOATS. The upwelling longwave and upwelling shortwave (ground temperature & albedo) was determined from the pyranometers and pyrgeometers on the gondola. The downwelling longwave (sky temperature) was determined from the ground station pyrgeometer. Accidental shadows cast on the groundstation by the field team were filtered out in post-processing. For both our flights, this sky temperature was cooler than the ground temperature.
Lastly, the direct shortwave (sunlight) was computed from a combination of all four pyranometers as none permanently face the morning sun, and the Lambert cosine effect on all sensors must accordingly be inverted. A sidewelling shortwave flux magnitude $E_{\text{sidewelling, solar}}$ was first computed as an envelope function from the side-mounted pyranometers, accounting for the motion in and out of the sunlight as the aerobot rotates. Subsequently, the upwelling flux from the bright ground was subtracted from this measured sideways signal with a viewfactor of 1/2 to provide a sun-only horizontal flux, and then direct solar magnitude was derived by removing the solar angle effect as:
@@
$=\sqrt{\left(E_{\text{sidewelling, solar}}-\frac{1}{2}E_{\text{upwelling, solar}}\right)^{2}+E_{\text{downwelling, solar}}^{2}}$ (28)
Given the early morning test flights with the sun close to the horizon and dim ground, however, the Lambert cosine effect has little influence. The resultant direct solar flux is nearly identical to measured sidewelling signal, but the distinction would be important for future longer duration flights closer to local noon.
All inputs could be supplied to FLOATS as functions of either altitude or time, depending on the phenomenon. We chose to parameterize the measured shortwave radiation and winds by time, as solar incidence angle is primarily time-dependent at low altitude, and the vertical wind gusts proved to be non-recurring disturbances. We decided to parameterize the measured atmospheric pressure and temperature by altitude instead of time (with high confidence for pressure and medium confidence on temperature; see Results Section 4), as these properties showed repetition on the aerobot ascent and descent - presumably given the long timescale of the atmosphere change in the morning compared to the balloon motion. Upwelling longwave radiation was nearly constant, but showed minor recurrence between ascent and descent and therefore was also parameterized by altitude. Downwelling longwave was measured at the groundstation so therefore parameterized by time.
@@ 4 Results
The main output comparisons for performance evaluation of the FLOATS model (namely aerobot altitude, SP balloon pressure, and gas and skin temperature) are shown in Figures 13, 13, and 14. Note that all plots zoom the Y-scale into the signal range in question to magnify the readability of data, and accordingly the zero value is not along the X-axis.
Aerobot altitude predictions from FLOATS (Fig.13) have both been tailored with free-lift values (in increments of 50 g) to best match the maximum altitude from the FLOATS simulation, as uncertainties of order 300 g were possible given our experience with the measurement technique in the groundwinds. For Flight 1, the over 4 m/s ground winds severely impacted our ability to accurately measure free lift, and tailored value of 100 g was significantly less than measured (350 g). In Flight 2 however, lower groundwinds meant our tailored value of 420 g was much closer to the measured value (320 g). We accordingly have much more confidence in altitude match of FLOATS in Flight 2. The phasing of rises and descents over the course of the flight matches well to the FLOATS simulation, meaning the instrumentation mostly captured the important short duration relevant inputs (winds/BCM actions etc). Accordingly, we believe the internal gas pressures were well measured and that the wind sensor was not shadowed by aerobot wake.
Fig. 13 plots predicted and observed SP absolute pressure values in both flights. We show the absolute pressure value of the gas within the SP balloon (absolute = gauge pressure ++ atmospheric pressure) as the gauge pressure is less diagnostic as it is sensitive to the precise altitude. We notice good agreement between the observed and simulated values, with approximately 2 kPa accumulated error over time ($\sim$ 2.5 $\%$). This error accumulation is dominated by the venting actions - the gas flow due to venting action is ZP pressure dependent (and therefore altitude dependent), so as FLOATS underestimates in altitude early in both flights it then underpredicts the vent actions. The SP gauge pressure rise rate due to the pump is captured better by the model, as it is a positive displacement pump with low sensitivity to altitude error.
We also note good agreement (Fig. 14) between the simulated and observed gas and skin temperatures with the ZP gas temperature. The temperature sensor measurements at the top and bottom of the aerobot are the most consistent with FLOATS predictions, whereas internal and side temperatures are less consistent. Typical thermal error between FLOATS predictions and temperature measurements is within $\pm 7^{\circ}$C, with better agreement during Flight 2 compared to Flight 1, and the most error is in the T3 sensors that cycle in and out of shadow rather than average into a single FLOATS node. The sudden temperature rise in FLOATS after activating the termination poppet is likely physical, as the ZP gas remaining after termination will adiabatically compress during rapid descent, but is not captured by the slow internal temperature logger that only logs every 5 minutes. In contrast, skin temperature sensors are sampled at 20 sps and averaged over 10 seconds to produce a time series at 0.1 sps do capture the rise. Future flights will emphasize better thermal access, especially within the ZP balloon to capture the thermally-driven buoyancy dynamics.
Figure 12: Simulated and observed altitudes and observed vertical velocity on (a) Flight 1 and (b) Flight 2. Green sections indicate venting, red sections indicate pumping, and blue indicates termination by the poppet. We assume a tailored 100 grams (Flight 1) and 420 grams (Flight 2) of free lift in FLOATS rather than measured values of 350 grams and 320 grams.
Figure 13: Simulated and observed SP balloon absolute pressure (not gauge pressure) for (a) Flight 1 and (b) Flight 2. As before, vent and pump actions are labeled in green and red respectively. Pump noise is an artifact of compression waves near the pressure transducer from the oscillating pump pistons.
Figure 14: Simulated and observed node temperatures. Note that the T2 node and SP node were not measured, and T3 was measured in two locations on opposite sides of the ZP envelope (shown by black and gray solid lines). Corrections from the post-flight calibration were applied (see appendix)
In our analysis, we found that the heat transfer model within FLOATS is the largest source of sensitivity in the outputs and likely drives our modeling error. Notably, the simulated aerobot altitude showcases the strongest sensitivity to the atmospheric temperature in flight. FLOATS cannot currently accommodate a time-carying atmosphere and only accepts the measured atmospheric temperature with altitude – thus variations in temperature at the same altitude during the approximately hour-long flights of the aerobot were averaged. This is a key limitation of the model, especially in the late morning as hot air near the ground begins to rise and mix, decreasing the air density. By contrast, FLOATS does simulate the rise in temperature of the ZP helium gas in sunlight. Accordingly, while FLOATS includes the increasing buoyancy effect from the warm helium it simultaneously ignores the decreasing buoyancy effect from the warm air.
@@
Future FLOATS modeling work will focus on these thermal excursions and requires longer duration flights for validation under many hours of variable solar forcing and atmospheric temperature change. Better capturing of the thermal time constant between the ZP gas and outside atmosphere (convective heat transfer rates), azimuthal nodes, and a time-varying atmospheric temperature model should improve the FLOATS prediction on how these variations impact the flight dynamics. Longer duration flights will also decrease the effect of free-lift measurement error, as uncertainty in the initial flight phase would get averaged over many hours.